OpenStudy (anonymous):

A circle is described by the equation x2 + y2 + 14x + 2y + 14 = 0. What are the coordinates for the center of the circle and the length of the radius? A) (-7, -1), 36 units B) (7, 1), 36 units C) (7, 1), 6 units D) (-7, -1), 6 units

OpenStudy (anonymous):

\[ x^2 + y^2 + 14x + 2y + 14 = 0\\ x^2 +14 x + 49 -49 +y^2+2y +1 -1 + 14=0\\ (x+7)^2 -49 + (y+1)^2 -1 + 14 =0 \] Can you finish it?

OpenStudy (anonymous):

\[ (x+7)^2 -49 + (y+1)^2 -1 + 14 =0\\ (x+7)^2 + (y+1)^2 =49+1 -14\\ (x+7)^2 + (y+1)^2 =36=6^2\\ \] What is the answer?

OpenStudy (anonymous):

Not really. Im not that great with math.

OpenStudy (anonymous):

The center is (-7,-1) and the radius is 6

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

YW